A boy went out into the barangay on his bike at 5kph. He rested two hrs. and returned home along the same road at the rate of 3 kph. If he was gone for 6 hrs, how far was the barangay from his home?
Answer by mananth(16949)

let the barangay distance be x km
riding speed = 5kph
riding time forward = x/5 hours
...
riding time return = x/3 hours
...
rest = 2 hours
..
x/5 + x/3 + 2 = 6
LCD = 30
(6x+10x+60)/30=6
16x+60 = 30*6
16x+60=180
16x=180-60
16x=120
x=7.5 km the distance to barangay
...
m.ananth@hotmail.ca